The sum of one-half of a number, one-ninth of the number, and one-twelfth of the number is twenty-five. What is the original number?
Answer by algebrahouse.com(1659)

The sum of one-half of a number, one-ninth of the number, and one-twelfth of the number is twenty-five. What is the original number?
Here's one way to do it:

x = the number

1/2(x) + 1/9(x) 1/12(x) = 25
36[1/2(x) + 1/9(x)+ 1/12(x) = 25] {multiplied entire equation by the LCD, 36}
36(1/2x) + 36(1/9x) + 36(1/12x) = 36(25) {multiplied each term by 36}
18x + 4x + 3x = 900 {multiplied each term by 36}
25x = 900 {combined like terms}
x = 36 {divided both sides by 36}
www.algebrahouse.com